# Moderate Adaptive Linear Units (MoLU)

### v5

@@ Abstract
We propose a new high-performance activation function, Moderate Adaptive Linear Units (MoLU), for the deep neural network. The MoLU, defined as : $f(x)=x\times\tanh(\alpha\times\exp(\beta\times x))$, is a simple, beautiful and powerful activation function that can be a good main activation function among hundreds of activation functions. Because the MoLU is made up of the elementary functions, not only it is a $C^{\infty}$-diffeomorphism (i.e. analytic over whole domains), but also it reduces the training time.
@@ II.1 MoLU
The MoLU is a simple, beautiful and powerful activation function that consists of a combination of hyperbolic tangent and exponential functions. The slope of the MoLU in the negative integer region make it possible to escape from the local minima. On the other hand, in the positive integer region, the slope of the MoLU is almost always unity, so that it allows us to find the global minimum of a loss function in a stable manner like ReLU. The MoLU is defined as below in Eq. (1).
$MoLU=x\times\tanh(\alpha\times\exp(\beta\times x))$ (1)
@@
Input GeLU Swish Mish ELU($\alpha=1$) MoLU($\alpha=2,\beta=2$)
-7 -2.33146835e-15 -6.37735836e-03 -6.38026341e-03 -9.99088118e-01 -1.16414021e-05
-6 -8.43964898e-11 -1.48357389e-02 -1.48540805e-02 -9.97521248e-01 -7.37305482e-05
-5 -2.29179620e-07 -3.34642546e-02 -3.35762377e-02 -9.93262053e-01 -4.53999296e-04
-4 -7.02459482e-05 -7.19448398e-02 -7.25917408e-02 -9.81684361e-01 -2.68370062e-03
-3 -3.63739208e-03 -1.42277620e-01 -1.45647461e-0 -9.50212932e-01 -1.48723912e-02
-2 -4.54023059e-02 -2.38405844e-01 -2.52501483e-01 -8.64664717e-01 -7.32298040e-02
-1 -1.58808009e-01 -2.68941421e-01 -3.03401461e-01 -6.32120559e-01 -2.64248689e-01
@@ II.2 Motivation
We realized that our activation function is not only showing a good performance for the accuracy but also converging to zero rapidly when updating a loss function during a test on some mathematical model and neural networks. We formulated our activation function in the following order. First, we used the hyperbolic tangent function as a basic framework, and then multiplied by the identity function to show the behavior of the identity function in the positive integer region. Lastly, we composited the exponential function to the hyperbolic tangent function to make it converge to zero asymptotically in the negative integer region. When we expanded the Mish[6] using a Taylor series, surprisingly, we happened to know that our activation function is related to the Mish.
${Mish}=$ $\ x\times\tanh(\log(1+e^{x}),\ \ e^{x}\leq 1$
$=$ $\ x\times\tanh(e^{x}-\frac{1}{2}e^{2x}+\frac{1}{3}e^{3x}-\cdots)$
$\approx$ $\ x\times\tanh(e^{x}),\ \ x<<1$
$=$ $\ {MoLU(\alpha=1,\beta=1)}$ (2)
As you can see Mish has a constraint since it has a logarithm function, but MoLu does not. However, this wonderful coincidence inspired us why our activation function works well at a faster rate.
@@ III Experiment
The robust property of MoLU is to approach rapidly to the value of minimum of a loss function without losing stability. This is a truly useful characteristic when training long time-series data by using NeuralODEs(Neural Ordinary Differential Equations). To prove the performance of MoLU, we conducted experiment on NeuralODEs, MNIST, and CIFAR10. In NeuralODEs, the differentiable activation functions are mainly used, so we compared MoLU with GeLU, Mish, SiLU, ELU, Tanh, and in case of the classification, compared it with ReLU, Leaky ReLU and Tanh. We used $\alpha=2,\beta=2$.
@@ III.1 NeuralODEs
Figure 4: Average of train losses and standard deviations. Mean value is the average of the minimum of the train loss for each three random seeds.
Table 2: NeuralODEs. Train losses and standard errors for some activation function
Train losses (×$\times 10^{-2}$)/Standard errors(×$\times 10^{-3}$)
MoLU GeLU Mish SiLU ELU Tanh
Train loss 2.02 2.85  3.69  1.77   4.47 7.03
Std. err. 2.74 11.57  21.02  1.17   5.41 5.16

### v6

@@ Abstract
We introduce the Moderate Adaptive Linear Units (MoLU), a novel activation function for deep neural networks, defined as: $f(x)=x\times\frac{1}{2}(1+\tanh(x))$. MoLU is a mathematically elegant and it exhibits outstanding performance in terms of accuracy, training speed, and computational efficiency, placing it as a versatile candidate among modern activation functions. Since MoLU is $C^{\infty}$-smooth (i.e., infinitely differentiable and analytic), it is expected to avoid the vanishing/exploding gradient problem and can be reliably applied across a wide range of domains, including LLMs, NeuralODEs, PINNs, and CNNs. Empirical evaluations demonstrate that MoLU achieves competitive convergence rates and higher accuracy compared to ReLU, Mish, and GeLU, while also reducing training time. As a result, MoLU stands out as a robust main activation function among the vast landscape of existing alternatives.
@@ II.1 MoLU
The MoLU is defined as below in Eq. (1).
$MoLU=x\times\frac{1}{2}(1+\tanh(x))$ (1)
@@
Input GeLU Swish($\beta=1$   ) Mish ELU($\alpha=1$) MoLU
-7 -2.33146835e-15 -6.37735836e-03 -6.38026341e-03 -9.99088118e-01 -5.82069619e-06
-6 -8.43964898e-11 -1.48357389e-02 -1.48540805e-02 -9.97521248e-01 -3.68650476e-05
-5 -2.29179620e-07 -3.34642546e-02 -3.35762377e-02 -9.93262053e-01 -2.26989344e-04
-4 -7.02459482e-05 -7.19448398e-02 -7.25917408e-02 -9.81684361e-01 -1.34140052e-03
-3 -3.63739208e-03 -1.42277620e-01 -1.45647461e-0 -9.50212932e-01 -7.41786947e-03
-2 -4.54023059e-02 -2.38405844e-01 -2.52501483e-01 -8.64664717e-01 -3.59724199e-02
-1 -1.58808009e-01 -2.68941421e-01 -3.03401461e-01 -6.32120559e-01 -1.19202922e-01
@@ II.2 Motivation
We realized that our activation function is not only showing a good performance for the accuracy but also converging to zero rapidly when updating a loss function during a test on some mathematical model and neural networks. We formulated our activation function in the following order. First, we used the hyperbolic tangent function as the foundational framework. To adapt its output range from (-1, 1) to (0, 1), we applied an affine transformation: adding 1 and dividing by 2. This scaled version, $\frac{1}{2}(1+tanh(x))$, ensures non-negative outputs while preserving smoothness.
Next, we multiplied this modified tanh function by the identity function $x$ to retain linear behavior in the positive region, analogous to the Swish activation function.
Finally, the inherent asymptotic convergence of tanh to -1 in the negative region ensures that the composite function $x\cdot\frac{1}{2}(1+tanh(x))$ naturally approaches zero as $x\rightarrow\infty$, eliminating the need for explicit exponential components.
@@ III Experiment
The robust property of MoLU is to approach rapidly to the value of minimum of a loss function without losing stability. This is a truly useful characteristic when training long time-series data by using NeuralODEs(Neural Ordinary Differential Equations). To prove the performance of MoLU, we conducted experiment on NeuralODEs, MNIST, and CIFAR10. In NeuralODEs, the differentiable activation functions are mainly used, so we compared MoLU with GeLU, Mish, SiLU, ELU, Tanh, and in case of the classification, compared it with ReLU, Leaky ReLU and Tanh.
@@ III.1 NeuralODEs
It is worth noting that our implementation of MoLU was defined manually as a custom Python function, without any low-level optimization. In contrast, the other activation functions were invoked directly from torch.nn.functional, which are internally optimized for performance within the PyTorch framework. Despite this disadvantage, MoLU not only achieved the highest accuracy but also exhibited the fastest training convergence among all tested activation functions.
Figure 4: Mean train losses with standard deviations. (Top) Total training times extracted from wandb. (Bottom)
Table 2: Summary of mean training losses with standard deviations and training time for NeuralODEs using various activation functions.
Mean train loss($10^{-2}$) / Standard error($10^{-3}$)    / Time (min)
 MoLU  GeLU  SiLU  Mish  ELU Tanh
Train loss 2.25 2.47  3.10  3.94   5.48 12.7
Standard dev.  3.13 2.10  8.67  7.01   11.1 7.80
Training time  18.6 21.7  19.1  19.3   18.8 20.2